LeeAnne’s bike has tires with 18-inch diameters. She rides her bike for 12 minutes, at an average speed of 10mi/hour. How many times do her tires rotate during this ride?
Answer by checkley77(12844)

12/60*10=1/5*10=2 MILES IS HER DISTANCE TRAVELED.
2*5,280=10,560 FT.
10,560/1.5=7,040 REVOLUTIONS OF THE TIRE FOR A 2 MILE RIDE.